# Phi-3 Technical Report: A Highly Capable Language Model Locally on Your Phone

### v2

@@ Abstract
We introduce phi-3-mini, a 3.8 billion parameter language model trained on 3.3 trillion tokens, whose overall performance, as measured by both academic benchmarks and internal testing, rivals that of models such as Mixtral 8x7B and GPT-3.5 (e.g., phi-3-mini achieves 69% on MMLU and 8.38 on MT-bench), despite being small enough to be deployed on a phone. The innovation lies entirely in our dataset for training, a scaled-up version of the one used for phi-2, composed of heavily filtered web data and synthetic data. The model is also further aligned for robustness, safety, and chat format. We also provide some initial parameter-scaling results with a 7B and 14B models trained for 4.8T tokens, called phi-3-small and phi-3-medium, both significantly more capable than phi-3-mini (e.g., respectively 75% and 78% on MMLU, and 8.7 and 8.9 on MT-bench).
@@ 1 Introduction
The striking progress of AI in the last few years can be largely attributed to major efforts throughout the world towards scaling-up to ever-larger models and datasets. Large Language Models (LLMs) have steadily increased in size from a mere billion parameters just five years ago (GPT-2 had 1.5 billion parameters [RWC+19]) to trillion parameters today. The impetus for this effort originates in the seemingly predictable improvement one obtains by training large models, the so-called scaling laws [KMH+20, HBM+22, MRB+23]. However these laws assume a “fixed” data source. This assumption is now significantly disrupted by the existence of frontier LLMs themselves, which allow us to interact with data in novel ways. In our previous works on the phi models [GZA+23, LBE+23, JBA+23] it was shown that a combination of LLM-based filtering of web data, and LLM-created synthetic data, enable performance in smaller language models that were typically seen only in much larger models. For example our previous model trained on this data recipe, phi-2 (2.7B parameters), matched the performance of models $25$ times larger trained on regular data. In this report we present a new model, phi-3-mini (3.8B parameters), trained for 3.3T tokens on larger and more advanced versions of the datasets used in phi-2. With its small size, phi-3-mini can easily be inferenced locally on a modern phone (see Figure 1), yet it achieves a quality that seems on-par with models such as Mixtral 8x7B [JSR+24] and GPT-3.5.
@@ 2 Technical Specifications
The phi-3-mini model is a transformer decoder architecture [VSP+17], with default context length $4K$. We also introduce a long context version via LongRope [DZZ+24] that extends the context length to $128K$, called phi-3-mini-128K.
@@
The phi-3-small model (7B parameters) leverages the tiktoken tokenizer (for better multilingual tokenization) with a vocabulary size of 100352 and has default context length $8K$. It follows the standard decoder architecture of a 7B model class, having $32$ layers and a hidden size of $4096$. To minimize KV cache footprint, the model also leverages a grouped-query attention, with $4$ queries sharing $1$ key. Moreover phi-3-small uses alternative layers of dense attention and a novel blocksparse attention to further optimize on KV cache savings while maintaining long context retrieval performance. An additional 10% multilingual data was also used for this model.
@@
We follow the sequence of works initiated in “Textbooks Are All You Need” [GZA+23], which utilize high quality training data to improve the performance of small language models and deviate from the standard scaling-laws. In this work we show that such method allows to reach the level of highly capable models such as GPT-3.5 or Mixtral with only 3.8B total parameters (while Mixtral has 45B total parameters for example). Our training data of consists of heavily filtered web data (according to the “educational level”) from various open internet sources, as well as synthetic LLM-generated data. Pre-training is performed in two disjoint and sequential phases; phase-1 comprises mostly of web sources aimed at teaching the model general knowledge and language understanding. Phase-2 merges even more heavily filtered webdata (a subset used in Phase-1) with some synthetic data that teach the model logical reasoning and various niche skills.
@@
Unlike prior works that train language models in either “compute optimal regime” [HBM+22] or “over-train regime”, we mainly focus on the quality of data for a given scale. [^2] Just like for “compute optimal regime”, we use the term “optimal” in an aspirational sense for “data optimal regime”. We are not implying that we actually found the provably “optimal” data mixture for a given scale. We try to calibrate the training data to be closer to the “data optimal” regime for small models. In particular, we filter the web data to contain the correct level of “knowledge” and keep more web pages that could potentially improve the “reasoning ability” for the model. As an example, the result of a game in premier league in a particular day might be good training data for frontier models, but we need to remove such information to leave more model capacity for “reasoning” for the mini size models. We compare our approach with Llama-2 in Figure 2.
@@
To test our data on larger size of models, we also trained phi-3-medium, a model with 14B parameters using the same tokenizer and architecture of phi-3-mini, and trained on the same data for slightly more epochs (4.8T tokens total as for phi-3-small). The model has 40 heads and 40 layers, with embedding dimension 5120. We observe that some benchmarks improve much less from 7B to 14B than they do from 3.8B to 7B, perhaps indicating that our data mixture needs further work to be in the “data optimal regime” for 14B parameters model. We are still actively investigating some of those benchmarks (including a regression on HumanEval), hence the numbers for phi-3-medium should be considered as a “preview”.
@@ 3 Academic benchmarks
On the next page we report the results for phi-3-mini on standard open-source benchmarks measuring the model’s reasoning ability (both common sense reasoning and logical reasoning). We compare to phi-2 [JBA+23], Mistral-7b-v0.1 [JSM+23], Mixtral-8x7b [JSR+24], Gemma 7B [TMH+24], Llama-3-instruct-8b [AI23], and GPT-3.5. All the reported numbers are produced with the exact same pipeline to ensure that the numbers are comparable. These numbers might differ from other published numbers due to slightly different choices in the evaluation. As is now standard, we use few-shot prompts to evaluate the models, at temperature 00. The prompts and number of shots are part of a Microsoft internal tool to evaluate language models, and in particular we did no optimization to the pipeline for the phi-3 models. [^3] For example, we found that using ## before the Question can lead to a noticeable improvement to phi-3-mini’s results across many benchmarks, but we did not do such changes in the prompts. The number of $k$–shot examples is listed per-benchmark. An example of a 2-shot prompt is described in Appendix A.
Phi-3-mini 3.8b Phi-3-small 7b (preview) Phi-3-medium 14b (preview) Phi-2 2.7b Mistral 7b Gemma 7b Llama-3-In 8b Mixtral 8x7b GPT-3.5 version 1106 MMLU (5-Shot) [HBK+21] 68.8 75.3 78.2 56.3 61.7 63.6 66.0 68.4 71.4 HellaSwag (5-Shot) [ZHB+19] 76.7 78.7 83.0 53.6 58.5 49.8 69.5 70.4 78.8 ANLI (7-Shot) [NWD+20] 52.8 55.0 58.7 42.5 47.1 48.7 54.8 55.2 58.1 GSM-8K (0-Shot; CoT) [CKB+21] 82.5 88.9 90.3 61.1 46.4 59.8 77.4 64.7 78.1 MedQA (2-Shot) [JPO+20] 53.8 58.2 69.4 40.9 49.6 50.0 58.9 62.2 63.4 AGIEval (0-Shot) [ZCG+23] 37.5 45.0 48.4 29.8 35.1 42.1 42.0 45.2 48.4 TriviaQA (5-Shot) [JCWZ17] 64.0 59.1 75.6 45.2 72.3 75.2 73.6 82.2 85.8 Arc-C (10-Shot) [CCE+18] 84.9 90.7 91.0 75.9 78.6 78.3 80.5 87.3 87.4 Arc-E (10-Shot) [CCE+18] 94.6 97.1 97.8 88.5 90.6 91.4 92.3 95.6 96.3 PIQA (5-Shot) [BZGC19] 84.2 87.8 87.7 60.2 77.7 78.1 77.1 86.0 86.6 SociQA (5-Shot) [BZGC19] 76.6 79.0 80.2 68.3 74.6 65.5 73.2 75.9 68.3 BigBench-Hard (0-Shot) [SRR+22, SSS+22] 71.7 75.0 81.3 59.4 57.3 59.6 68.9 69.7 68.32 WinoGrande (5-Shot) [SLBBC19] 70.8 82.5 81.4 54.7 54.2 55.6 58.0 62.0 68.8 OpenBookQA (10-Shot) [MCKS18] 83.2 88.4 87.2 73.6 79.8 78.6 81.6 85.8 86.0 BoolQ (0-Shot) [CLC+19] 77.2 82.9 86.6 72.2 66.0 78.3 77.6 79.1 CommonSenseQA (10-Shot) [THLB19] 80.2 80.3 82.6 69.3 72.6 76.2 73.6 78.1 79.6 TruthfulQA (10-Shot) [LHE22] 65.0 68.7 75.7 52.1 53.0 62.0 60.1 85.8 HumanEval (0-Shot) [CTJ+21] 59.1 59.1 55.5 47.0 28.0 34.1 60.4 37.8 62.2 MBPP (3-Shot) [AON+21] 70.0 71.4 74.5 60.6 50.8 51.5 65.3 60.2 77.8 Average 71.2 74.9 78.2 61.0 62.0 68.0 69.9 75.3 GPQA (2-Shot; CoT) [RHS+23] 32.8 34.3 29.0 MT Bench (2 round ave.) [ZCS+23] 8.38 8.70 8.91 8.35
@@ 4 Safety
Table 1 shows the results of in-house RAI benchmarks for phi-3-mini-4k and phi-3-mini-128k compared to phi-2 [JBA+23], Mistral-7b-v0.1 [JSM+23], Gemma 7b [TMH+24], and Llama-3-instruct-8b [AI23]. This benchmark utilized GPT-4 to simulate multi-turn conversations in five different categories and to evaluate the model responses. Ungroundedness between 0 (fully grounded) and 4 (not grounded) measures if the information in a response is based on a given prompt. In other categories, responses were evaluated in terms of the severity of harmfulness from 0 (no harm) to 7 (extreme harm) and the defect rates (DR-$x$) were computed as the percentage of samples with the severity score being greater than or equal to $x$.
Phi-3-Mini-4k 3.8b Phi-3-Mini-128k 3.8b Phi-2 2.7b Mistral 7b Gemma 7b Llama-3-In 8b Ungroundedness 0.603 0.637 1.481 0.935 0.679 0.328 Intellectual Property (DR-1) 23.95% 21.50% 24.00% 56.20% 38.33% 37.30% Harmful Content Continuation (DR-3) 0.75% 1.08% 2.93% 2.58% 1.28% 1.30% Harmful Content Summarization (DR-3) 10.00% 10.20% 14.35% 22.33% 10.33% 8.20% Jailbreak (DR-1) 12.29% 12.57% 15.00% 15.57% 11.43% 13.00%
Table 1: Comparison of Microsoft internal multi-turn conversation RAI benchmark results of phi-3-mini and other models. Note that a lower value indicates a better performance for all metrics in the table.
@@ 5 Weakness
Despite our diligent RAI efforts, as with most LLMs, there remains challenges around factual inaccuracies (or hallucinations), reproduction or amplification of biases, inappropriate content generation, and safety issues. The use of carefully curated training data, and targeted post-training, and improvements from red-teaming insights significantly mitigates these issues across all dimensions. However, there is significant work ahead to fully address these challenges.
Figure 4: Left: phi-3-mini’s completion without search. Right: phi-3-mini’s completion with search, using the default HuggingFace Chat-UI search ability.

### v3

@@ Abstract
We introduce phi-3-mini, a 3.8 billion parameter language model trained on 3.3 trillion tokens, whose overall performance, as measured by both academic benchmarks and internal testing, rivals that of models such as Mixtral 8x7B and GPT-3.5 (e.g., phi-3-mini achieves 69% on MMLU and 8.38 on MT-bench), despite being small enough to be deployed on a phone. The innovation lies entirely in our dataset for training, a scaled-up version of the one used for phi-2, composed of heavily filtered publicly available web data and synthetic data. The model is also further aligned for robustness, safety, and chat format. We also provide some initial parameter-scaling results with a 7B and 14B models trained for 4.8T tokens, called phi-3-small and phi-3-medium, both significantly more capable than phi-3-mini (e.g., respectively 75% and 78% on MMLU, and 8.7 and 8.9 on MT-bench). Moreover, we also introduce phi-3-vision, a 4.2 billion parameter model based on phi-3-mini with strong reasoning capabilities for image and text prompts.
@@ 1 Introduction
The striking progress of AI in the last few years can be largely attributed to major efforts throughout the world towards scaling-up to ever-larger models and datasets. Large Language Models (LLMs) have steadily increased in size from a mere billion parameters just five years ago (GPT-2 had 1.5 billion parameters [RWC+19]) to trillion parameters today. The impetus for this effort originates in the seemingly predictable improvement one obtains by training large models, the so-called scaling laws [KMH+20, HBM+22, MRB+23]. However these laws assume a “fixed” data source. This assumption is now significantly disrupted by the existence of frontier LLMs themselves, which allow us to interact with data in novel ways. In our previous works on the phi models [GZA+23, LBE+23, JBA+23] it was shown that a combination of LLM-based filtering of publicly available web data, and LLM-created synthetic data, enable performance in smaller language models that were typically seen only in much larger models. For example our previous model trained on this data recipe, phi-2 (2.7B parameters), matched the performance of models $25$ times larger trained on regular data. In this report we present a new model, phi-3-mini (3.8B parameters), trained for 3.3T tokens on larger and more advanced versions of the datasets used in phi-2. With its small size, phi-3-mini can easily be inferenced locally on a modern phone (see Figure 2), yet it achieves a quality that seems on-par with models such as Mixtral 8x7B [JSR+24] and GPT-3.5.
@@ 2 Technical Specifications
The phi-3-mini model is a transformer decoder architecture [VSP+17], with default context length $4K$. We also introduce a long context version via LongRope [DZZ+24a] that extends the context length to $128K$, called phi-3-mini-128K.
@@
The phi-3-small model (7B parameters) leverages the tiktoken tokenizer (for better multilingual tokenization) with a vocabulary size of 100352222We remove unused tokens from the vocabulary. and has default context length $8192$. It follows the standard decoder architecture of a 7B model class, having $32$ heads, $32$ layers and a hidden size of $4096$. We switched from GELU activation to GEGLU and used Maximal Update Parametrization (muP) [yang2022tensor] to tune hyperparameters on a small proxy model and transfer them to the target 7B model. Those helped ensure better performance and training stability. Also, the model leverages a grouped-query attention, with $4$ queries sharing $1$ key. To optimize the training and inference speed, we design a novel blocksparse attention module. For each attention head, the blocksparse attention enforces different sparsity patterns over KV cache. This ensures that all tokens are attended to on different heads for the given choice of sparsity. As illustrated in Figure 1, the context is then efficiently divided and conquered among attention heads, with significant KV cache reduction. To achieve actual deployment speed-up from the blocksparse design, we implemented highly efficient, yet flexible kernels for both training and inference. For training, we build a triton kernel based on Flash Attention [DFE+22]. For inference, we implemented a kernel for the prefilling phase and extended the paged attention kernel in vLLM for the decoding phase [KLZ+23]. Lastly, in phi-3-small architecture, we alternate dense attention layers and blocksparse attention layers to optimize KV cache savings while maintaining long context retrieval performance. An additional 10% multilingual data was also used for this model.
Figure 1: Toy illustration of the blocksparse attention in phi-3-small with 2 local blocks and vertical stride of 3. The table shows the Keys/values a query token in block 8 attended to. Blue=local blocks, orange=remote/vertical blocks, gray=blocks skipped.
@@
We follow the sequence of works initiated in “Textbooks Are All You Need” [GZA+23], which utilize high quality training data to improve the performance of small language models and deviate from the standard scaling-laws. In this work we show that such method allows to reach the level of highly capable models such as GPT-3.5 or Mixtral with only 3.8B total parameters (while Mixtral has 45B total parameters for example). Our training data of consists of heavily filtered publicly available web data (according to the “educational level”) from various open internet sources, as well as synthetic LLM-generated data. Pre-training is performed in two disjoint and sequential phases; phase-1 comprises mostly of web sources aimed at teaching the model general knowledge and language understanding. Phase-2 merges even more heavily filtered webdata (a subset used in Phase-1) with some synthetic data that teach the model logical reasoning and various niche skills.
@@
Unlike prior works that train language models in either “compute optimal regime” [HBM+22] or “over-train regime”, we mainly focus on the quality of data for a given scale. [^3] Just like for “compute optimal regime”, we use the term “optimal” in an aspirational sense for “data optimal regime”. We are not implying that we actually found the provably “optimal” data mixture for a given scale. We try to calibrate the training data to be closer to the “data optimal” regime for small models. In particular, we filter the publicly available web data to contain the correct level of “knowledge” and keep more web pages that could potentially improve the “reasoning ability” for the model. As an example, the result of a game in premier league in a particular day might be good training data for frontier models, but we need to remove such information to leave more model capacity for “reasoning” for the mini size models. We compare our approach with Llama-2 in Figure 3.
@@
To test our data on larger size of models, we also trained phi-3-medium, a model with 14B parameters using the same tokenizer and architecture of phi-3-mini, and trained on the same data for slightly more epochs (4.8T tokens total as for phi-3-small. The model has 40 heads and 40 layers, with embedding dimension 5120. We observe that some benchmarks improve much less from 7B to 14B than they do from 3.8B to 7B, perhaps indicating that our data mixture needs further work to be in the “data optimal regime” for 14B parameters model.
@@ 3 Academic benchmarks
On the next page we report the results for phi-3-mini on standard open-source benchmarks measuring the model’s reasoning ability (both common sense reasoning and logical reasoning). We compare to phi-2 [JBA+23], Mistral-7b-v0.1 [JSM+23], Mixtral-8x7b [JSR+24], Gemma 7B [TMH+24], Llama-3-instruct-8b [AI], and GPT-3.5. All the reported numbers are produced with the exact same pipeline to ensure that the numbers are comparable. These numbers might differ from other published numbers due to slightly different choices in the evaluation. As is now standard, we use few-shot prompts to evaluate the models, at temperature 00. The prompts and number of shots are part of a Microsoft internal tool to evaluate language models, and in particular we did no optimization to the pipeline for the phi-3 models. [^4] For example, we found that using ## before the Question can lead to a noticeable improvement to phi-3-mini’s results across many benchmarks, but we did not do such changes in the prompts. The number of $k$–shot examples is listed per-benchmark. An example of a 2-shot prompt is described in Appendix A.
Phi-3-mini 3.8b Phi-3-small 7b Phi-3-medium 14b Phi-2 2.7b Mistral 7b Gemma 7b Llama-3-In 8b Mixtral 8x7b GPT-3.5 version 1106 MMLU (5-Shot) [HBK+21] 68.8 75.7 78.0 56.3 61.7 63.6 66.5 70.5 71.4 HellaSwag (5-Shot) [ZHB+19] 76.7 77.0 82.4 53.6 58.5 49.8 71.1 70.4 78.8 ANLI (7-Shot) [NWD+20] 52.8 58.1 55.8 42.5 47.1 48.7 57.3 55.2 58.1 GSM-8K (8-Shot; CoT) [CKB+21] 82.5 89.6 91.0 61.1 46.4 59.8 77.4 64.7 78.1 MedQA (2-Shot) [JPO+20] 53.8 65.4 69.9 40.9 50.0 49.6 60.5 62.2 63.4 AGIEval (0-Shot) [ZCG+23] 37.5 45.1 50.2 29.8 35.1 42.1 42.0 45.2 48.4 TriviaQA (5-Shot) [JCWZ17] 64.0 58.1 73.9 45.2 75.2 72.3 67.7 82.2 85.8 Arc-C (10-Shot) [CCE+18] 84.9 90.7 91.6 75.9 78.6 78.3 82.8 87.3 87.4 Arc-E (10-Shot) [CCE+18] 94.6 97.0 97.7 88.5 90.6 91.4 93.4 95.6 96.3 PIQA (5-Shot) [BZGC19] 84.2 86.9 87.9 60.2 77.7 78.1 75.7 86.0 86.6 SociQA (5-Shot) [BZGC19] 76.6 79.2 80.2 68.3 74.6 65.5 73.9 75.9 68.3 BigBench-Hard (3-Shot; CoT) [SRR+22, SSS+22] 71.7 79.1 81.4 59.4 57.3 59.6 51.5 69.7 68.32 WinoGrande (5-Shot) [SLBBC19] 70.8 81.5 81.5 54.7 54.2 55.6 65.0 62.0 68.8 OpenBookQA (10-Shot) [MCKS18] 83.2 88.0 87.4 73.6 79.8 78.6 82.6 85.8 86.0 BoolQ (2-Shot) [CLC+19] 77.2 84.8 86.5 72.2 66.0 80.9 77.6 79.1 CommonSenseQA (10-Shot) [THLB19] 80.2 80.0 82.8 69.3 72.6 76.2 79.0 78.1 79.6 TruthfulQA (10-Shot; MC2) [LHE22] 65.0 70.2 75.1 53.0 52.1 63.2 60.1 85.8 HumanEval (0-Shot) [CTJ+21] 58.5 61.0 62.2 59.0 28.0 34.1 60.4 37.8 62.2 MBPP (3-Shot) [AON+21] 70.0 71.7 75.2 60.6 50.8 51.5 67.7 60.2 77.8 Average 71.2 75.7 78.5 61.2 61.7 69.4 69.8 74.3 GPQA (2-Shot; CoT) [RHS+23] 32.8 34.3 29.0 MT Bench (2 round ave.) [ZCS+23] 8.38 8.70 8.91 8.35
@@ 4 Safety
The safety alignment of phi-3-small and phi-3-medium was conducted by undergoing the same red-teaming process, utilizing identical datasets, and incorporating a slightly larger number of samples. Table 1 shows the results of in-house RAI benchmarks [MHJ+23] for phi-3 models compared to phi-2 [JBA+23], Mistral-7b-v0.1 [JSM+23], Gemma 7b [TMH+24], and Llama-3-instruct-8b [AI]. This benchmark utilized GPT-4 to simulate multi-turn conversations in five different categories and to evaluate the model responses. Ungroundedness between 0 (fully grounded) and 4 (not grounded) measures if the information in a response is based on a given prompt. In other categories, responses were evaluated in terms of the severity of harmfulness from 0 (no harm) to 7 (extreme harm) and the defect rates (DR-$x$) were computed as the percentage of samples with the severity score being greater than or equal to $x$.
Phi-3-mini 3.8b Phi-3-small 7b Phi-3-medium 14b Phi-2 2.7b Mistral 7b Gemma 7b Llama-3-In 8b Ungroundedness 0.603 0.299 0.213 1.481 0.935 0.679 0.328 Third Party Harm (DR-1) 0.240 0.253 0.251 0.240 0.562 0.383 0.373 Harmful Content Continuation (DR-3) 0.007 0.003 0.010 0.029 0.026 0.013 0.013 Harmful Content Summarization (DR-3) 0.100 0.110 0.112 0.144 0.223 0.103 0.082 Jailbreak (DR-1) 0.123 0.107 0.111 0.150 0.156 0.114 0.130
Table 1: Comparison of Microsoft internal multi-turn conversation RAI benchmark results of phi-3 models and other models. Note that a lower value indicates a better performance for all metrics in the table.
@@ 5 Weakness
Despite our diligent RAI efforts, as with most LLMs, there remains challenges around factual inaccuracies (or hallucinations), reproduction or amplification of biases, inappropriate content generation, and safety issues. The use of carefully curated training data, and targeted post-training, and improvements from red-teaming insights significantly mitigates these issues across all dimensions. However, there is significant work ahead to fully address these challenges, and downstream use of the models should be evaluated for the specific use cases and safety considerations for that context.
Figure 5: Left: phi-3-mini’s completion without search. Right: phi-3-mini’s completion with search, using the default HuggingFace Chat-UI search ability. For reference, the 2026 Winter Olympic Games are scheduled to be held in Milano and Cortina in Italy, while the 2022 and 2018 Winter Olympic Games were held in Beijing, China and PyeongChang, Korea, respectively. Without the search results, the response is incorrect, while with the web search, not only does the response become accurate, but also gets more specific with suggestions.
6 Phi-3-Vision
Figure 6: The demo case shows Phi-3-Vision’s capability in natural image understanding and reasoning.
6.1 Technical Specifications
Architecture
The Phi-3-Vision (4.2B parameters) is a multimodal model designed to process an image and a textual prompt as inputs, and subsequently generate textual outputs. This model is composed of two primary components: an image encoder, i.e., CLIP ViT-L/14 [RKH+21] and a transformer decoder, i.e., phi-3-mini-128K-instruct. The visual tokens, once extracted by the image encoder, are then combined with text tokens in an interleaved way (no particular order for image and text tokens). To accommodate high-resolution images and various aspect ratios, a dynamic cropping strategy [DZZ+24b] is utilized to split the input image into a 2d array of blocks, where the tokens of the blocks are concatenated to represent the whole image.
Pre-training
The Phi-3-Vision model undergoes a pre-training phase using a diverse dataset, which consists of a combination of interleaved image-text documents (e.g.,  [LST+24]), image-text pairs from FLD-5B  [XWX+24], synthetic data derived from Optical Character Recognition (OCR) of PDF files, datasets for chart/table comprehension, and text-only data. The objective of predicting the next token is employed specifically on text tokens, while any loss associated with image tokens is disregarded during this phase. The pre-training process involves a total of $0.5T$ tokens that encompass both visual and text elements. During the pre-training phase, the maximum image resolution is capped at $1344\times 1344$ as the majority of the training images are smaller than this resolution.
Post-training.
The Phi-3-Vision model contains two post-training stages: supervised finetuning (SFT) and direct preference optimization (DPO). For SFT, we leveraged text SFT dataset, public multimodal instruct tuning datasets along with large-scale multimodal instruct tuning datasets that we built ourselves, covering diverse domains and tasks such as general natural image understanding, chart/table/diagram understanding/reasoning, PowerPoint understanding, and model safety. The multimodal SFT data has about a total of 15B tokens. For DPO we mainly use a text DPO dataset and a relatively smaller-scale multimodal DPO dataset. For these two stages, we jointly train multimodal tasks and text-only tasks so that the model can achieve multi-modal reasoning while maintaining language capabilities as much as possible.
6.2 Academic benchmarks
We report in Table 2 the evaluation results of Phi-3-Vision on nine open-source academic benchmarks. These benchmarks evaluate reasoning and perceptual capabilities on visual and text inputs and can be grouped in three categories: Science, Charts, and Generic knowledge. We compare Phi-3-Vision with the following baselines: MM1-3B-Chat [MGF+24], MM1-7B-Chat [MGF+24], Llava-1.6 Vicuna 7B [LLLL23], Llava-1.6 Llama3-8B [LLL+24], Qwen-VL-Chat [BBY+23], Claude 3 Haiku [Ant24], Gemini 1.0 Pro V [TAB+23], and GPT-4V-Turbo. Our performance quality assessment setup used the same evaluation pipeline for all the baselines to ensure a fair comparison, with the exception of MM1-3B-Chat. We just copied and pasted their published numbers since the model is not publicly available.
Our evaluation setup aimed to mimic scenarios where regular users interact with a multi-modal model, i.e., users who are not experts in prompt engineering or know special techniques that can improve performance. For this reason, we adopted the evaluation setting used in Llava-1.5 [LLLL23]. In this setup, the prompts include instructions to select a single letter corresponding to an answer from a list of given options, or answer with a single word or phrase. In our prompts, we did not use specific tokens for multiple-choice questions. Moreover, we did not scale or pre-process any image in our benchmarking system. We placed the images as the first item in the prompts, except on the MMMU dataset where the prompts interleave the images anywhere in the question or the answers. Lastly, our evaluation setup only considered a 0-shot format. Because of these evaluation parameters, our reported numbers can differ from the published numbers of the considered baselines.
Phi-3-Vision 4.2b MM1-3B-Chat 3.6b [MGF+24] MM1-7B-Chat 7.6b [MGF+24] LLaVA-1.6 Vicuna-7b [LLLL23] LLaVA-Next LLama3-8b [LLL+24] Qwen-VL-Chat 9.6b [BBY+23] Claude 3 haiku  [Ant24] Gemini 1.0 Pro V  [TAB+23] GPT-4V-Turbo turbo-2024-04-09 MMMU (val) [YNZ+23] 40.4 33.9 37.0 34.2 36.4 39.0 40.7 42.0 55.5 ScienceQA (test) [LMX+22] 90.8 69.4 72.6 70.6 73.7 67.2 72.0 79.7 75.7 MathVista (testmini) [LBX+24] 44.5 32.0 35.9 31.5 34.8 29.4 33.2 35.0 47.5 Inter-GPS (test) [LGJ+21] 38.1 - - 20.5 24.6 22.3 32.1 28.6 41.0 MMBench (dev-en) [LDZ+24] 80.5 75.9 79.0 76.3 79.4 75.8 62.4 80.0 86.1 POPE (test) [LDZ+23] 85.8 87.4 86.6 87.2 87.0 82.6 74.4 84.2 83.7 AI2D (test) [KSK+16] 76.7 - - 63.1 66.9 59.8 60.3 62.8 74.7 ChartQA (test) [MLT+22] 81.4 - - 55.0 65.8 50.9 59.3 58.0 62.3 TextVQA (test) [SNS+19] 70.9 71.9 72.8 64.6 55.7 59.4 62.7 64.7 68.1
Table 2: Comparison results on public MLLM benchmarks. All the reported numbers are produced with the exact same pipeline to ensure that the numbers are comparable except for MM1-3B-Chat [MGF+24] and MM1-7B-Chat [MGF+24], which are not publicly available. We adopted the evaluation setting used in Llava-1.5 [LLLL23], without any specific prompt or pre-processing image for all results. These numbers might differ from other published numbers due to slightly different prompts.
6.3 Safety
To ensure the integration of Phi-3-Vision aligns with Microsoft’s Responsible AI (RAI) principles, we involved safety post-training in both Supervised Fine-Tuning (SFT) stage and Direct Preference Optimization (DPO) stage. In creating the safety training datasets, we utilized not only the text-only RAI datasets, but also a variety of in-house Multi-Modal (MM) RAI datasets that cover various harm categories identified in both public and internal MM RAI benchmarks. For the purpose of RAI evaluation, we performed a rigorous quantitative assessment on both public and internal benchmarks, this was done in conjunction with a human evaluation conducted by Microsoft’s internal red team.
Phi-3-Vision 3.8b+0.3b Phi-3-Vision w/o safety 3.8b+0.3b Llava-1.6 Vicuna 7b+0.3b Qwen-VL-Chat 7.7b+1.9b GPT4-V N/A Internal (private) 8.30 7.06 5.44 7.27 8.55 RTVLM (public) 4.64 3.56 3.86 4.78 6.81 VLGuard (public) 9.12 4.66 5.62 8.33 8.90
Table 3: Comparison results on public and private multi-modal RAI benchmarks. Note that all metrics in the table are [0,10] and a higher value indicates a better performance.
In Table 3, we present the evaluation outcomes of Phi-3-Vision on three MM RAI benchmarks: one internal and two public benchmarks (specifically, RTVLM [LLY+24] and VLGuard [ZBY+24]). We juxtapose these results with those of other open-source models such as Llava-1.5 [LLLL23], Llava-1.6 [LLL+24], Qwen-VL-Chat [BBY+23], and GPT4-V[Ope23]. The results clearly indicate that safety post-training notably enhances the RAI performance of Phi-3-Vision across all RAI benchmarks. In Figure 7, we further breakdown the performance across different RAI categories of the VLGuard and Internal benchmarks, demonstrating that safety post-training can aid Phi-3-Vision in improving RAI performance in nearly all categories.
Figure 7: Comparison of categorized RAI performance of Phi-3-Vision with and without the safety post-training on the VLGuard (left) and Internal (right) benchmark, respectively. It clearly indicates that safety post-training can enhance the RAI performance across nearly all the RAI categories.
6.4 Weakness
Regarding the multi-modal LLM capabilities of our Phi-3-Vision, it performs admirably across various fields. However, we have identified certain limitations, particularly with questions necessitating high-level reasoning abilities. Additionally, the model has been observed to occasionally generate ungrounded outputs, making it potentially unreliable in sensitive areas, such as finance. To mitigate these issues, we will incorporate more reasoning-focused and hallucination-related DPO data into post-training in the future.
From a responsible AI standpoint, whilst safety post-training has made significant strides, our Phi-3-Vision occasionally fails to refrain from answering harmful or sensitive inquiries. Examples of such occasions include deciphering particular types of captcha and describing scam images containing disinformation or hallucination. We find that this issue partly arises from the capabilities, such as OCR, acquired during the training process with normal instruct tuning datasets, which can be regarded as the trade-off between helpfulness and harmlessness. Moving forward, we need to further explore this area to achieve a better balance.